OpenStudy (anonymous):

solve using the elimination method. 5x-y=5, 3x+y=11, please help me to do this

OpenStudy (kropot72):

Add the two equations to eliminate y giving: 8x = 16 x = 2 Can you work out the value of y by substituting for x = 2 ?

OpenStudy (anonymous):

wow, i didnt konw you could just add the two up like that, I think my teacher showed us something way different

OpenStudy (kropot72):

Substituting for x in 3x + y = 11 gives: (3 * 2) + y = 11 6 + y = 11 y = 11 - 6 = 5

OpenStudy (anonymous):

your awesome, thanks

OpenStudy (anonymous):

keep an eye out for me if I need more help please, only three more pages to do and your help is making it easier than i thought

OpenStudy (kropot72):

You're welcome :) BTW when the equations are added, the like terms on each side are added. Equal expressions are added to each side. Subtraction can also be used to eliminate one of the variables. Experience will show whether addition or subtraction will minimise the working out.

OpenStudy (anonymous):

I posted another one i need help with, i would have asked you here but if i repost it i can give you another medal. some people really like them, i actually get asked for them